OpenStudy (anonymous):

x^3+x^2+x/-x dividing a polynomial by a monomial. Will this work if so will the answer be -x^3-x^2-x

jimthompson5910 (jim_thompson5910):

you did the sign changes correctly but you are forgetting that x^3 divided by x = x^(3-1) = x^2

OpenStudy (anonymous):

I knew i forgot something. How would you solve it?

jimthompson5910 (jim_thompson5910):

x^3/x = ??? x^2/x = ??? x/x = ???

OpenStudy (anonymous):

Thanks.

jimthompson5910 (jim_thompson5910):

yw